# Coherent Zero-Shot Visual Instruction Generation

### v1

@@ 1 Introduction
Directly inputting the instructions to the text-to-image models incurs several issues, as shown by the examples in Figure 2. First, most instructions explain the procedure of actions, while the text-to-image model expects the description of the image content. As shown in step 2 of the “Making hot chocolate” example, the container of the milk becomes a glass bowl instead of the steel pot as mentioned in step 1 due to the lack of information in the instruction. This necessitates an approach to bridging the gap between instructional texts and the conditioning text used for image generation. Formally, given the procedures at the current and previous steps, one needs to infer the proper image content at the current step. To this end, we propose an instruction re-captioning strategy [3, 67] to convert the instructional texts into actions and states using large language models (LLMs). We show that combining the action and state as the condition significantly enhances the quality and relevance of the generated illustrations.
@@
In addition, the objects’ identity may alter arbitrarily across different steps. As found in the “planting a houseplan” example, the shape and texture of the pot in the first step look different from the one in steps 3 and 4. This poses a common challenge encountered when using text-to-image models to generate multiple images - there often lacks coherency across the generated illustrations. Recent studies have made progress in maintaining consistency for human portraits [10, 21, 29], where the identity features can be derived from the models trained on extensive human-centric datasets. However, our problem cannot benefit from the same idea since the identity features of the general object categories, such as kitchen utensils, are unavailable. Therefore, we refer to the general methods to achieve consistency, such as feature sharing and injection [69, 71].
@@
Although these methods improve the generation consistency, we find that they also induce the “over-consistent” problem in visual instruction generation. Specifically, many instructions involve changed objects among different steps. For instance, a recipe may contain a raw ingredient that is chopped, seasoned, and presented in entirely different states throughout the process. This inherent variability complicates the problem of maintaining object consistency and makes it unsuitable to use the vanilla feature-sharing strategy. To this end, we propose an adaptive feature-sharing method with finer-grained constraints. First, similar to the previous method, we adopt a local region constraint to enforce sharing to happen on certain pixels. However, we find that localizing the objects using attention maps becomes less reliable when the base model architecture differs from the UNet in Stable Diffusion [46]. Instead, we utilize large-scale segmentation models to produce the masks [35, 32, 38, 53]. Second, we apply a global constraint on the feature sharing scale between each pair of steps based on their similarity. We exploit the reasoning capabilities of the LLMs to develop a similarity matrix that characterizes such state similarity. We show that our adaptive feature-sharing method enables the variability of objects across instructional steps while still preserving object identity.
@@
As showcased in Figure 1, our proposed method can generate high-quality, consistent visual representations based on the instructional texts. We also perform quantitative experiments with various evaluation metrics to ablate individual components of our method on the text-image alignment and consistently. However, we notice that the image similarity metric only emphasizes the consistency aspect and conflicts with the goal of achieving variation across different steps. To this end, we propose a framework to evaluate the visual instruction generation quality using large-scale visual language models. We will release the code and full evaluation suite for reproducible research. The contributions of our work can be summarized as follows:
@@
    We propose an illustration re-captioning strategy, which greatly improves the generation quality and relevance.
@@ 2 Related work
Improving Consistency in Image Generation . Generating consistent images has been an important sub-problem in different tasks, including video generation [19, 31, 73, 79], multi-view image generation [60, 39, 70], and character generation [72, 10, 21, 29, 69]. Comparing with these tasks, we focus on improving the consistency of the shared objects in different steps of the visual instructions. Unlike video or multi-view image generation, we don’t enforce hard-constraint geometrical or temporal consistency. Different from the character generation that only cares about human identity, we need to deal with general objects that don’t have ID feature extractors as those for humans [72, 10, 21, 29].
@@ 3 Method
Given a set of instructions, we harness a pre-trained text-to-image diffusion model to generate the visual illustrations. As shown in Fig. 3, our approach contains two major stages. First, to fill the distribution gap between the instructions and image descriptions, we perform in-context planning with LLMs to re-caption the instructions. Second, we propose an adaptive feature-sharing method for dynamic, coherent image generation given the re-captioned instructions. In both stages, we use off-the-shelf pre-trained models without any extra training.
Figure 3: Our framework for zero-shot instruction visualization. Our framework operates in two distinct phases. In the first phase, we use an LLM (e.g., the GPT-4 model) to generate the scene state after each step in the list of instructions. The generated scene state helps guide the image generation in the next stage. We also ask the LLM to generate the similarity between states. This matrix, with each row, indicating the visual similarity of a current visual step to others, guides the generation process. For example, to achieve high state similarity, we wish to maintain consistency as much as possible across the two steps. A low state similarity indicates the performed action changes the scene state substantially. In such cases, blindly encouraging consistency across steps may hurt the quality of the visualized instruction image. In the second phase, we utilize a shared attention layer—replacing the standard model—to allow queries from one image to access keys and values from others within the same instruction set. We enhance this sharing mechanism by applying standard attention masking, controlled by the similarity matrix, to tune the interaction between visual elements finely.
@@ 3.1 Preliminaries
Self-attention layer is essential in $D$ for integrating global information across the entire image. It redistributes the features from each spatial location to similar regions. Suppose that $\mathbf{x}\in\mathbb{R}^{w\times h\times d}$ denote the input feature map of some self-attention layer, where $w$, $h$, and $d$ are the width, height, and dimension. For simplicity, let $P=h\times w$. By applying linear mappings to the feature map $\mathbf{x}$ to obtain the key $K\in\mathbb{R}^{P\times d_{k}}$, value $V\in\mathbb{R}^{P\times d_{v}}$, and query $Q\in\mathbb{R}^{P\times d_{k}}$ where $d_{k}$ is dimension of key $K$ and query $Q$, $d_{v}$ is dimension of value $V$, the self-attention map $A^{t}$ at step $t$ is generated by :
@@
Attention maps mechanism calculates the similarity between query (Q) and key (K), which determine how much attention each value (V) receives.
@@ 3.2 Re-captioning instructions as descriptive texts
An immediate solution to this issue is to concatenate the adjacent instructions. However, this may lead to conflicted information as the instructions express different actions. (It can seen in the second row of the Figure  9). Therefore, we propose to leverage LLMs’ conversational understanding capabilities to first re-caption the instruction into detailed input texts.
@@ 3.3 Dynamic consistent image generation
Our re-captioning technique ensures the description continuity in text prompts, while we still need to generate consistent content shared across instructions. Different from previous methods like video or story generation, which require absolute consistency on the entire scene or the object region, instructional generation requires a more nuanced consistency. As we discussed in the previous sections, some visual instruction steps may demand maintaining object consistency, while others may involve significant transformations of the objects or background according to the procedure. To this end, we propose an adaptive KV sharing method based on local region and state similarity constraints.
KV sharing with local region constraint. Similar to Consistory [68], visual instruction generation only cares about consistency within certain regions. We identify these consistent objects in the instructions during our in-content planning phase. With our re-captioned instructions, the images generated realize most information about the instructions, like the layout, except for the object consistency. Therefore, we propose to use off-the-shelf segmentation models to generate the object masks [38, 53]. We also explored using cross-attention maps for producing the mask as in Consistory. However, we find that the link between text and image can be weak and noisy in the state-of-the-art text-to-image models [47] as shown in Figure 4. Instead, we show that existing segmentation models work well with generated images and are not overfitted to a single model architecture.
@@
KV sharing with state similarity constraint. To manage the dynamic scenario of generating consistent objects with variations, we generalize the KV sharing in a continuous manner. Instead of only sharing or not sharing KV features among different regions, we regulate the scaling of sharing with a state similarity matrix. This matrix is designed to control the degree of similarity between each pair of instructional steps. By adjusting its values, one can tailor the visual output to either highlight consistency or emphasize transformation, depending on the instructions. To automatically infer this matrix from the instructions, we provide thoughtful instruction and in-context examples to LLMs. Let $W\in[0,1]^{N\times N}$ be the state similarity matrix, where $N$ is the number of steps. $W_{i,j}$ represents the overall similarity between the $i$-th and the $j$-th steps.
@@
Here, $A_{i}^{+}$ represents the adjusted attention weights, where the flow of information in self-attention is scaled inversely by the values in $S_{i}^{+}$. When $S_{j}$ approaches zeros (indicating no similarity between steps $i$-th and $j$-th states), the sharing of information between those specific steps is nullified. The magnitude of information sharing between the $i$-th and $j$-th steps is substantial when the values of $S_{j}$ are high, which encourages more consistent regions to be generated. This method allows us to precisely control the trade-off between consistency and variation across different steps in the visual instruction generation process, ensuring that each generated step is appropriately aligned with both the instruction and the other images in the sequence.
@@ 4.1 Experiment setup
Evaluation Following the previous studies, we utilize CLIP-Score  [51] to measure the text-image alignment, and Dreamsim  [16] and L2_Dinov2  [8] to evaluate the consistency. However, the tasks exhibited in the instructions span a wide range of categories. Each of these may have a distinct goal that goes beyond simple object consistency. Understanding the quality of the visual illustration demands nuanced reasoning—for instance, while some scenarios necessitate maintaining object consistency, others require deliberate changes in the object’s state, such as cutting or decorating. Therefore, we find that conventional metrics are often insufficient to evaluate such a complex task. Instead, we turn to recent VLMs like GPT-4V and Gemini-Pro 1.5 which show great visual understanding and reasoning capacity. We mainly evaluate in the following aspects:
@@
To apply these metrics, we use carefully designed instructions and in-context examples to query vision-language models. At each time, we provide a pair of generated images, one by our method and the other by the baseline method. The VLM is requested to pick the image better at the above aspects.
@@
Implementation details. Our method includes two main components. We use GPT-4 API for in-context planning. We provide all the in-context examples and prompts in the appendix. However, we note that users can achieve similar quality using ChatGPT3 or 4. For dynamic, consistent image generation, we use Stable Cascade  [47] as our text-to-image model, which has three stages: A, B, and C. Based on the report and exploration, we find that only stage C forms the image based on the text condition (stage B uses text as a condition, but in this stage, the text condition barely affects the generation results). Thus, we apply our adaptive KV sharing methodin stage C. Specifically, we apply it in the first 15 steps of a total of 20.
@@ 4.2 Quantitative results
We evaluate our method and baseline approaches using vision-language models and show results in Figure 5 and Figure 6. We quantitatively compare using different conditioning texts as input in Figure 5, demonstrating that our re-captioning approach achieves overall better results than baseline methods that rely solely on instructions for image generation. Our method shows consistently improved behavior across all four aspects. We then quantitatively compare different feature sharing methods in Figure  6. We find that our adaptive KV sharing methods improve almost all metrics using Gemini and GPT-4V.
@@
Figure 6: Evaluation of different consistency methods using LLM, including Gemini and GPT-4. Using both masks and weights achieves overall the best performances among all the choices.
@@
Table 1: Ablation study of different design choices of the text prompts. We find that using instructions only provides the best text image alignment while concatenating with previous instructions or states improves the coherency.
@@ 4.3 Qualitative results
Figure 8: Visual comparisons of consistency across steps. Here, we use the proposed prompting approach and focus on validating the various ways of sharing keys and values in attention layers for consistent image generation. All methods take the same input text prompt. KV: sharing across early steps; KV local: sharing controlled by masks; KV global: sharing controlled by the proposed state similarity matrix; Ours: sharing controlled by both masks and the state similarity matrix. For example, in the omelette, nal̈ively sharing the key and value across steps cause content leaking (e.g., the pan in step 3 and 4 looks like the bowl in step 1 and 2). With weight control, the method can maintain consistency, avoiding leaking feature while respecting the action. Adding mask control helps improve textual alignment. (e.g, the actions are more align with text)
@@
Figure  9 illustrates the impact of various design choices for prompts on the coherence and alignment of the generated instructions. When only the action from a single step is used as input, the generated images lack coherence due to unawareness of the context. For instance, in “cooking Hot Chocolate”, the pot is only mentioned in step 1. As a result, steps 2 and 3 do not have such information that the pot is in the stove and contains the hot milk. Therefore, the model fails to maintain the context; the pot is not in the stove and does not contain the milk anymore in steps 2 and 3. In scenarios where current and previous actions are concatenated, the model generally can understand the context better. However, such models tend to prioritize the most recent action at the cost of accurately generating earlier ones. Concatenating action and state significantly enhances context understanding, leading to more accurate image generation based on prior steps. For instance, this method ensures the generated image accurately places the pot on the stove in the case of a pot and cooking. In another example involving cookies, the approach allows for generating images of cookies with chocolate chips, ensuring the action aligns closely with the input text.
@@
In Figure  10, we compare our method with GenHowTo [65], a pretrained model for generating actions and states from real images. We use prompts from the GenHowTo test dataset, which belongs to the same categories as their training data. Note that our method is training-free and does not require a real input image as input. Ours produces results that are comparable in terms of textual alignment, consistency, and even better visual quality. This is because our approach leverages a state-of-the-art text-to-image model. Unlike GenHowTo  [65], which requires re-training the model for each new base text-to-image model and may compromise the generation quality due to the training data quality.

### v2

@@ 1 Introduction
Directly inputting the instructions to the text-to-image models incurs several issues, as shown by the examples in Figure 2. First, most instructions explain the procedure of actions, while the text-to-image model expects the description of the image content. As shown in step 2 of the “Making hot chocolate” example, the milk container becomes a glass bowl instead of the steel pot as mentioned in step 1 due to the lack of information in the instruction. This necessitates an approach to bridging the gap between instructional texts and the conditioning text used for image generation. Given the procedures at the current and previous steps, one needs to infer the proper image content at the current step. To this end, we propose an instruction re-captioning strategy [3, 69] to convert the instructional texts into actions and states using large language models (LLMs). We show that combining the action and state as the condition significantly enhances the quality and relevance of the generated illustrations.
@@
In addition, the objects’ identities may alter arbitrarily across different steps. For example, in the ”planting a houseplan” example, the shape and texture of the pot in the first step are different from those in steps 3 and 4. This poses a common challenge when using text-to-image models to generate multiple images - there is often a lack of coherency across the generated illustrations. Recent studies have made progress in maintaining consistency for human portraits [10, 21, 29], where the identity features can be derived from the models trained on extensive human-centric datasets. However, our problem cannot benefit from the same idea since the identity features of the general object categories, such as kitchen utensils, are unavailable. Therefore, we refer to the general methods to achieve consistency, such as feature sharing and injection [71, 73].
@@
Although these methods improve the generation consistency, they also induce the “over-consistent” problem in visual instruction generation. Specifically, many instructions involve changed objects among different steps. For instance, a recipe may contain a raw ingredient that is chopped, seasoned, and presented in entirely different states throughout the process. This inherent variability complicates the problem of maintaining object consistency and makes it unsuitable to use the vanilla feature-sharing strategy. To this end, we propose an adaptive feature-sharing method with finer-grained constraints. First, similar to the previous method, we adopt a local region constraint to enforce sharing to happen on certain pixels. However, localizing the objects using attention maps becomes less reliable when the base model architecture differs from the UNet in Stable Diffusion [48]. Instead, we utilize large-scale segmentation models to produce the masks [37, 34, 40, 55, 30, 31]. Second, we apply a global constraint on the feature-sharing scale between each pair of steps based on their similarity. We exploit the reasoning capabilities of the LLMs to develop a similarity matrix that characterizes such state similarity. We show that our adaptive feature-sharing method enables the variability of objects across instructional steps while preserving object identity.
@@
As showcased in Figure 1, our proposed method can generate high-quality, consistent visual representations based on instructional texts. We also perform quantitative experiments with various evaluation metrics to ablate individual components of our method on the text-image alignment and consistency. However, we notice that the image similarity metric only emphasizes the consistency aspect and conflicts to achieve variation across different steps. To this end, we propose a framework to evaluate the visual instruction generation quality using large-scale visual language models. We will release the code and full evaluation suite for reproducible research. The contributions of our work can be summarized as follows:
@@
    We propose an illustration re-captioning strategy, greatly improving the generation quality and relevance.
@@ 2 Related work
Improving Consistency in Image Generation . Generating consistent images has been an important sub-problem in different tasks, including video generation [19, 33, 75, 81], multi-view image generation [62, 41, 72], and character generation [74, 10, 21, 29, 71]. Compared with these tasks, we focus on improving the consistency of the shared objects in different steps of the visual instructions. Unlike video or multi-view image generation, we don’t enforce hard-constraint geometrical or temporal consistency. Different from the character generation that only cares about human identity, we need to deal with general objects that don’t have ID feature extractors as those for humans [74, 10, 21, 29].
@@ 3 Method
Given a set of instructions, we harness a pre-trained text-to-image diffusion model to generate the visual illustrations. As shown in Fig. 3, our approach contains two major stages. First, to fill the distribution gap between the instructions and image descriptions, we perform in-context planning with LLMs to re-caption the instructions. Second, given the re-captioned instructions, we propose an adaptive feature-sharing method for dynamic, coherent image generation. In both stages, we use off-the-shelf pre-trained models without any extra training.
Figure 3: Our framework for zero-shot instruction visualization. Our framework operates in two distinct phases. In the first phase, we use an LLM (e.g., the GPT-4 model) to generate the scene state after each step in the list of instructions. The generated scene state helps guide the image generation in the next stage. We also ask the LLM to generate the similarity between states. This matrix, with each row indicating the visual similarity of a current visual step to others, guides the generation process. For example, to achieve high state similarity, we wish to maintain consistency as much as possible across the two steps. A low state similarity indicates the performed action changes the scene state substantially. In such cases, blindly encouraging consistency across steps may hurt the quality of the visualized instruction image. In the second phase, we utilize a shared attention layer—replacing the standard model—to allow queries from one image to access keys and values from others within the same instruction set. We enhance this sharing mechanism by applying standard attention masking, controlled by the similarity matrix, to finely tune the interaction between visual elements.
@@ 3.1 Preliminaries
Self-attention layer is essential in $D$ for integrating global information across the entire image. It redistributes the features from each spatial location to similar regions. Suppose that $\mathbf{x}\in\mathbb{R}^{w\times h\times d}$ denote the input feature map of some self-attention layer, where $w$, $h$, and $d$ are the width, height, and dimension. Let $P=h\times w$ for simplicity. By applying linear mappings to the feature map $\mathbf{x}$ to obtain the key $K\in\mathbb{R}^{P\times d_{k}}$, value $V\in\mathbb{R}^{P\times d_{v}}$, and query $Q\in\mathbb{R}^{P\times d_{k}}$ where $d_{k}$ is dimension of key $K$ and query $Q$, $d_{v}$ is dimension of value $V$, the self-attention map $A^{t}$ at step $t$ is generated by :
@@
The attention maps mechanism calculates the similarity between query (Q) and key (K), which determines how much attention each value (V) receives.
@@ 3.2 Re-captioning instructions as descriptive texts
An immediate solution to this issue is to concatenate the adjacent instructions. However, this may lead to conflicting information as the instructions express different actions. (It can seen in the second row of Figure 9). Therefore, we propose leveraging LLMs’ conversational understanding capabilities to re-caption the instruction into detailed input texts.
@@ 3.3 Dynamic consistent image generation
Our re-captioning technique ensures the description continuity in text prompts, while we still need to generate consistent content shared across instructions. Unlike previous methods like video or story generation, which require absolute consistency on the entire scene or the object region, instructional generation requires a more nuanced consistency. As discussed in the previous sections, some visual instruction steps may demand maintaining object consistency, while others may involve significant transformations of the objects or background according to the procedure. To this end, we propose an adaptive KV-sharing method based on local region and state similarity constraints.
KV sharing with local region constraint. Similar to Consistory [70], visual instruction generation only cares about consistency within certain regions. We identify these consistent objects in the instructions during our in-content planning phase. With our re-captioned instructions, the images generated realize most of the information about the instructions, like the layout, except for the object consistency. Therefore, we propose to use off-the-shelf segmentation models to generate the object masks [40, 55]. We also explored using cross-attention maps to produce the mask, as in Consistory. However, the link between text and image can be weak and noisy in the state-of-the-art text-to-image models [49] as shown in Figure 4. Instead, we show that existing segmentation models work well with generated images and are not overfitted to a single model architecture.
@@
KV sharing with state similarity constraint. To manage the dynamic scenario of generating consistent objects with variations, we continuously generalize the KV sharing. Instead of only sharing or not sharing KV features among different regions, we regulate the scaling of sharing with a state similarity matrix. This matrix controls the degree of similarity between each pair of instructional steps. By adjusting its values, one can tailor the visual output to highlight consistency or emphasize transformation, depending on the instructions. To automatically infer this matrix from the instructions, we provide thoughtful instruction and in-context examples to LLMs. Let $W\in[0,1]^{N\times N}$ be the state similarity matrix, where $N$ is the number of steps. $W_{i,j}$ represents the overall similarity between the $i$-th and the $j$-th steps.
@@
Here, $A_{i}^{+}$ represents the adjusted attention weights, where the flow of information in self-attention is scaled inversely by the values in $S_{i}^{+}$. When $S_{j}$ approaches zeros (indicating no similarity between steps $i$-th and $j$-th states), the sharing of information between those specific steps is nullified. The magnitude of information sharing between the $i$-th and $j$-th steps is substantial when the values of $S_{j}$ are high, encouraging more consistent regions to be generated. This method allows us to precisely control the trade-off between consistency and variation across different steps in the visual instruction generation process, ensuring that each generated step is appropriately aligned with the instruction and the other images in the sequence.
@@ 4.1 Experiment setup
Evaluation Following the previous studies, we utilize CLIP-Score  [53] to measure the text-image alignment, and Dreamsim  [16] and L2_Dinov2  [8] to evaluate the consistency. However, the tasks in the instructions span a wide range of categories. Each of these may have a distinct goal beyond simple object consistency. Understanding the quality of the visual illustration demands nuanced reasoning—for instance. At the same time, some scenarios necessitate maintaining object consistency, and others require deliberate changes in the object’s state, such as cutting or decorating. Therefore, conventional metrics often do not evaluate such a complex task. Instead, we turn to recent VLMs like GPT-4V and Gemini-Pro 1.5, which show great visual understanding and reasoning capacity. We mainly evaluate in the following aspects:
@@
To apply these metrics, we use carefully designed instructions and in-context examples to query vision-language models. We provide a pair of generated images, one by our method and the other by the baseline method, each time. The VLM is requested to pick the image better in the above aspects.
@@
Implementation details. Our method includes two main components. We use GPT-4 API for in-context planning. We provide all the in-context examples and prompts in the appendix. However, we note that users can achieve similar quality using ChatGPT3 or 4. For dynamic, consistent image generation, we use Stable Cascade  [49] as our text-to-image model, which has three stages: A, B, and C. Based on the report and exploration, we find that only stage C forms the image based on the text condition (stage B uses text as a condition, but in this stage, the text condition barely affects the generation results). Thus, we apply our adaptive KV sharing method in stage C. Specifically, we apply it in the first 15 steps, which total 20.
@@ 4.2 Quantitative results
We evaluate our method and baseline approaches using vision-language models and show results in Figure 5 and Figure 6. We quantitatively compare using different conditioning texts as input in Figure 5, demonstrating that our re-captioning approach achieves overall better results than baseline methods relying solely on image generation instructions. Our method shows consistently improved behavior across all four aspects. We then quantitatively compare different feature-sharing methods in Figure  6. Our adaptive KV sharing methods improve almost all metrics using Gemini and GPT-4V.
@@
Figure 6: Evaluation of different consistency methods using LLM, including Gemini and GPT-4. Using both masks and weights achieves the best performances among all the choices overall.
@@
Table 1: Ablation study of different design choices of the text prompts. Using instructions only provides the best text image alignment while concatenating with previous instructions or states improves the coherency.
@@ 4.3 Qualitative results
Figure 8: Visual comparisons of consistency across steps. Here, we use the proposed prompting approach and focus on validating the various ways of sharing keys and values in attention layers for consistent image generation. All methods take the same input text prompt. KV: sharing across early steps; KV local: sharing controlled by masks; KV global: sharing controlled by the proposed state similarity matrix; Ours: sharing controlled by both masks and the state similarity matrix. For example, in the omelet, nal̈ively sharing the key and value across steps cause content leaking (e.g., the pan in steps 3 and 4 looks like the bowl in steps 1 and 2). With weight control, the method can maintain consistency, avoiding leaking feature while respecting the action. Adding mask control helps improve textual alignment. (e.g, the actions are more align with text)
@@
Figure  9 illustrates the impact of various design choices for prompts on the coherence and alignment of the generated instructions. When only the action from a single step is used as input, the generated images lack coherence due to unawareness of the context. For instance, in “cooking Hot Chocolate”, the pot is only mentioned in step 1. As a result, steps 2 and 3 do not know that the pot is in the stove and contains the hot milk. Therefore, the model fails to maintain the context; the pot is not in the stove and does not contain the milk anymore in steps 2 and 3. In scenarios where current and previous actions are concatenated, the model can better understand the context. However, such models tend to prioritize the most recent action at the cost of accurately generating earlier ones. Concatenating action and state significantly enhances context understanding, leading to more accurate image generation based on prior steps. For instance, this method ensures the generated image accurately places the pot on the stove in the case of a pot and cooking. In another example involving cookies, the approach allows for generating images of cookies with chocolate chips, ensuring the action aligns closely with the input text.
@@
In Figure  10, we compare our method with GenHowTo [67], a pretrained model for generating actions and states from real images. We use prompts from the GenHowTo test dataset, which belongs to the same categories as their training data. Note that our method is training-free and does not require a real input image as input. Ours produces comparable results regarding textual alignment, consistency, and even better visual quality. This is because our approach leverages a state-of-the-art text-to-image model. Unlike GenHowTo  [67], which requires re-training the model for each new base text-to-image model and may compromise the generation quality due to the training data quality.